OpenStudy (anonymous):

PLZ IMMEDIATE HELP!! The perimeter of a rectangle is 36ft. The length is twice the width. Find length and width of da rectabgle.. WILL FAN & MEDAL I PROMISE!

OpenStudy (ipwnbunnies):

Do you know the general formula for perimeter?

OpenStudy (anonymous):

no I forgot -_-

OpenStudy (ipwnbunnies):

.-. Perimeter is the length of the boundary of a rectangle. So, Perimeter = 2*length + 2*width

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so.... can I just plz know the length and width first... I need to answer and explain, so answer and explain for me ( explain in shortest way possible plz )

OpenStudy (ipwnbunnies):

Since length is twice the width, you can substitute length like this. P = 2L + 2W Since, L = 2W P = 2(2W) + 2W Solve for W, that'll give you the width. And you can find the length. c:

OpenStudy (anonymous):

i donz understand.. this...

OpenStudy (anonymous):

plz i just need answer now...

OpenStudy (anonymous):

or else.... bad stuff happens

OpenStudy (ipwnbunnies):

Remember, I said Perimeter = 2*length + 2*width Length = 2*width, according to the problem. Sorry, I'm here to help. Not give answers.

OpenStudy (anonymous):

ok fine -_-

OpenStudy (anonymous):

sooo.......

OpenStudy (ipwnbunnies):

P = 2*length + 2*width Replace 'length' with '2*width' P = 2*(2*width) + 2*width P = 4*width + 2*width P = 6*width You're given the perimeter. Solve for width.

OpenStudy (anonymous):

6???

OpenStudy (ipwnbunnies):

Mhm, the width is 6. And the problem tells you that the length is twice the width. Voila.

OpenStudy (anonymous):

6 and 12?? i did it??? YAAAAAAAAYAYYAYAYAYAYAYAYAY thank u!!!!

OpenStudy (ipwnbunnies):